OpenStudy (anonymous):

write in point-slope form the equation of the line through each pair of points. (1/2,2/3) and (-3/5, 5/3)

OpenStudy (anonymous):

slope is -1/2??

OpenStudy (anonymous):

First pair 3-2/2-1=1 so y2-y1=1(x2-x)

OpenStudy (anonymous):

whoa whoa whoa. what??? is my slope right???

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

y2-y1/x2-x1= Slope

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

what??? i dont even know how you got 1 as you slope

OpenStudy (anonymous):

what first pair?

OpenStudy (anonymous):

(1/2,2/3)

OpenStudy (anonymous):

you want me to find the slope of that???

OpenStudy (anonymous):

i can't find the slope with one pair.

OpenStudy (anonymous):

write in point slope form.

OpenStudy (anonymous):

i see then (y2-y1)=m(x2-x1)

OpenStudy (anonymous):

i know. i got my slope as -2/1

OpenStudy (anonymous):

-2/8

OpenStudy (anonymous):

that is insane... how on earth did you get that.

OpenStudy (anonymous):

y-5= -2/8(x+3

OpenStudy (anonymous):

3-5/5-(-3)

OpenStudy (anonymous):

like this vid? http://www.youtube.com/watch?v=qtiCRNDd-zI hahaha

OpenStudy (anonymous):

those coordinates arent right.

OpenStudy (anonymous):

after you get the slopes just plug then in http://www.purplemath.com/modules/strtlneq2.htm